OpenStudy (anonymous):

The inverse of negative Pi over three? Not entirely sure how to approach this question. Please help?

OpenStudy (loser66):

to me, it's 3/pi

OpenStudy (anonymous):

Try that with sqrt(3): y/x=sqrt(3) y=x*sqrt(3) Solve Pythagoras but substituting y: x² + (x*sqrt(3))² = 1 x² + x² * 3 = 1 4x² = 1 x² = 1/4 => x=1/2, y=sqrt(3)/2 You should then RECOGNIZE the "square root 3 over 2" value as a side of a 30-60-90 triangle in a unit circle. Since y>x, x is positive and arctan is also positive, it is a 60 degree or pi/3 angle

OpenStudy (anonymous):

Thank you.